OpenStudy (anonymous):

What is the perimeter of a rectangle with length six and one over two inches and width 3 inches? 9 inches nine and one over two inches 16 inches 19 inches @dg2

OpenStudy (anonymous):

perimeter of rectangle=2(l+b)

OpenStudy (anonymous):

l-length b-breadth substitute in that formula

OpenStudy (anonymous):

2(6.5+3) 19

OpenStudy (anonymous):

length=13/2 breadth =3 so perimeter =2(13/2+3)=19

OpenStudy (anonymous):

xanny give dg a medal for me please.

OpenStudy (anonymous):

yup